OpenStudy (anonymous):

four pencils and five pens costs $2.00, three pencils and four pens costs $1.58 Find the cost of a pencil and the cost of a pen

OpenStudy (amistre64):

x = pencil; y = pen 4x + 5y = 2 y=2-4x//5 ................... 3x + 4y = 1.58 3x + 4(2-4x//5) = 1.58 3x+ 8/5 -16x/5 = 1.58 15x/5 -16x/5 = 1.58 - 8/5 -x/5 = 1.58 - 8/5 x = -5(1.58 - 8/5)

OpenStudy (amistre64):

pencils = -7.90+8 = .10

OpenStudy (amistre64):

y=2-4x//5 y=2-4(.10)//5 y = 10/5 - .40/5 = 9.60/5 = 1.92 pens = 1.92

OpenStudy (amistre64):

or not lol

OpenStudy (amistre64):

pencil = .40 pen = .32 had the right eqautions, but forgot how to divide....